### v2

Abstract
Causal weighted quantile treatment effects (WQTE) are a useful complement to standard causal contrasts that focus on the mean when interest lies at the tails of the counterfactual distribution. To-date, however, methods for estimation and inference regarding causal WQTEs have assumed complete data on all relevant factors. In most practical settings, however, data will be missing or incomplete data, particularly when the data are not collected for research purposes, as is the case for electronic health records and disease registries. Furthermore, such data sources may be particularly susceptible to the outcome data being missing-not-at-random (MNAR). In this paper, we consider the use of double-sampling, through which the otherwise missing data are ascertained on a sub-sample of study units, as a strategy to mitigate bias due to MNAR data in the estimation of causal WQTEs. With the additional data in-hand, we present identifying conditions that do not require assumptions regarding missingness in the original data. We then propose a novel inverse-probability weighted estimator and derive its asymptotic properties, both pointwise at specific quantiles and uniformly across a range of quantiles over some compact subset of $(0,1)$, allowing the propensity score and double-sampling probabilities to be estimated. For practical inference, we develop a bootstrap method that can be used for both pointwise and uniform inference. A simulation study is conducted to examine the finite sample performance of the proposed estimators. The proposed method is illustrated with data from an EHR-based study examining the relative effects of two bariatric surgery procedures on BMI loss at 3 years post-surgery.
keywords:
Bootstrap, Causal inference, Quantile regression, Quantile regression process, Heterogeneous treatment effect, Missing not at random, Uniform inference, Unspecified missing data mechanism.
1 Introduction
Modern causal inference methods conceptualize causation via a comparison of counterfactuals or potential outcomes across interventions or treatment strategies (Hernán and Robins, 2020). Because one cannot hope to learn about causation for individuals, it is standard to focus on contrasts of some summary statistic of the treatment-specific counterfactual distributions. The most common such contrast, arguably, is the average treatment effect (ATE), although related contrasts have been considered including the average treatment effect on the treated (e.g., Rubin, 1977; Imbens, 2004), the survivor average treatment effect (e.g., Robins, 1986; Tchetgen Tchetgen, 2014), and average effects of modified treatment policies (e.g., Muñoz and van der Laan, 2012; Haneuse and Rotnitzky, 2013). That each of these contrasts focuses on the mean of the counterfactual distribution, in one way or another, is natural since the mean is standard measure of centrality. In some settings, however, particularly when low-probability but high-impact events/outcomes are anticipated, understanding the causal effect of a given treatment beyond the ‘center’, including at the tails of the counterfactual distributions, may be more relevant to decision-making (Sun et al., 2021; Meyerson et al., 2023). In such settings, quantile treatment effects (QTEs), defined as the difference in quantiles of counterfactual distributions across treatment options, may be a useful compliment to standard contrasts based on the mean (e.g., Andrews et al., 2016; Schiele and Schmitz, 2016; Autor et al., 2017; Hemilä et al., 2022).
Towards estimation of causal QTEs, there is an emerging literature most of which focuses on inverse-probability weighting (IPW) using the propensity score (Rosenbaum and Rubin, 1983; Firpo, 2007; Sun et al., 2021). That methods analogous to the g-formula or doubly-robust methods have not been developed is, in part, due to Jensen’s inequality precluding the characterization of population QTEs as a function of covariate-conditional QTEs. Throughout this literature, and in contrast to the literature on related but distinct quantile regression methods (Yi and He, 2009; Yuan and Yin, 2010; Wei et al., 2012; Zhao et al., 2013; Sherwood et al., 2013; Han et al., 2019; Yu et al., 2023), it has been implicitly assumed that complete data is available on the outcome and all relevant covariates. Missing or incomplete data, however, is a well-known and widespread practical challenge in almost all research settings where human subjects are involved. This is particularly the case when the data at-hand were not collected for research purposes as in the case with electronic health records (EHRs), claims-based databases and registries (Haneuse and Shortreed, 2017). In studies of the long-term effects of bariatric surgery, for example, data on relevant outcomes (e.g., weight/BMI, hypertension, type 2 diabetes and mortality) are subject to a range of mechanisms that result in missing data, including: disenrollment from a health plan; that a patient doesn’t have a relevant health encounter at a time point of interest; and, that a measurement was not taken even though standards of care would indicate that it should have been (Haneuse and Daniels, 2016; Peskoe et al., 2021).
When it comes to performing analyses based on data that are incomplete, however, analysts must engage in two related tasks: (i) clarifying and invoking assumptions necessary for valid estimation and inference; and, (ii) employing some statistical methodology that acknowledges the missingness, as appropriate. Regarding the former, it is typical that the missing-at-random (MAR) assumption is invoked (Little and Rubin, 2019); intuitively, MAR corresponds to probability of missingness depending only on observed data. In some settings, however, MAR may not hold, in that missingness depends, in part at least, on quantities that are unobserved. This scenario is generally referred to as missing-not-at-random (MNAR) or informative missingness. Unfortunately, in the absence of additional information, MNAR cannot be empirically verified. In turn, much of the literature has focused on sensitivity analysis methods (e.g., Robins et al., 2000) and estimating bounds on the parameters of interest (e.g., Manski, 1990). An alternative approach, which we pursue in this work, is to engage in additional data collection efforts, specifically to (re)contact a subset of subjects with missing data to obtain complete data. With this data in-hand, otherwise untestable assumptions can be evaluated (at least partially) and/or incorporated into analyses. Such a strategy, sometimes referred to as double-sampling is well-established as a means to mitigate confounding bias (Breslow, 1996; Schildcrout et al., 2020), bias due to measurement error (Lynn et al., 2012; Bound et al., 2001) and ecological bias (Haneuse and Wakefield, 2007, 2008). However, with few exceptions (Frangakis and Rubin, 2001; An et al., 2009; Qian et al., 2020; Coppock et al., 2017; Levis et al., 2023), double-sampling has not been considered routinely as a means to mitigate bias due to potentially MNAR data.
With this collective backdrop, this paper proposes the use of double-sampling as a means to mitigate bias due to potentially informative missing data for a continuous outcome when interest lies in estimating contrasts in a class of weighted quantile treatment effects (WQTEs). These estimands, introduced by Sun et al. (2021), facilitate comparisons of potential outcomes over various selected target populations, such as among the treated. Towards this, we first put forth conditions sufficient for the identification of WQTEs in observational settings with missing outcomes, when the double-sampling design is pursued. Based on this, we propose an IPW approach to estimate WQTEs, labelled IPW-WQTE. We derive the asymptotic distribution of the proposed estimator with estimated propensity scores and missing mechanism at any fixed quantile level (i.e., pointwise inference), and establish uniform inference of the entire IPW-WQTEs process across all quantiles over a compact subset of $(0,1)$. Finally, we propose a bootstrap resampling-based method can be used for either pointwise or uniform inference. The performance of the proposed approach and estimator is evaluated via a simulation study, with proofs of the main theorems provided in the Supplementary Material. The proposed method is then illustrated with data from an EHR-based study examining the relative effects of two bariatric surgery procedures on BMI loss at 3 years post-surgery.
2 Weighted quantile treatment effects
In this section, we introduce notation, then briefly review the definition of WQTEs, as well as identification and estimation details in the absence of missing data.
2.1 Notation
Let $Y\in\mathbbm{R}$ denote the continuous outcome of interest and $Z\in\mathcal{Z}=\{0,1\}$ a binary treatment indicator (e.g., $Z$=$0/1$=control/treatment). The methods proposed in this paper are derived under binary treatment for simplicity, but can be generalized to categorical and continuous treatment using the generalized propensity score (e.g., Imbens, 2000; Sun et al., 2021). Let $X\in\mathcal{X}\subset\mathbbm{R}^{p}$ denote a $p$-dimensional vector of covariates that are sufficient for the control of confounding of the $Z\sim Y$ association. We refer to $(Y,Z,X^{\top})$ as the ‘complete data’ and conceive of it as arising from some joint distribution $P_{c}$ from which one could, in principle, derive, among other quantities: $F_{Y}(y)=\textrm{pr}(Y\leq y)$, the marginal cumulative distribution function of $Y$; and, $F_{Y\mid Z,X}(y)$ and $f_{Y\mid Z,X}(y)$ the conditional cumulative distribution and density functions of $Y$ given $Z$ and $X$, respectively.
Towards embedding this work within the causal inference paradigm, let $Y(z)$ denote the potential outcome corresponding to treatment level $Z=z$ for $z\in\mathcal{Z}$, and $F_{Y(z)}(y)$ and $F_{Y(z)\mid X}(y)$ denote the corresponding marginal and conditional cumulative distribution functions, respectively.
2.2 Weighted quantile treatment effects
Define the $g$-weighted cumulative distribution function for $Y(z)$,
$F^{g}_{Y(z)}(y)=\frac{E\left\{g(X)F_{Y(z)\mid X}(y\mid X)\right\}}{E\left\{g(X)\right\}},\quad z\in\{0,1\},$ (1)
where $g$ is a user-specified function of $X$ for which there exists $c_{g}>0$ such that $0<g(x)<c_{g}$ for all $x\in\mathcal{X}$. We then define the $\tau$th WTQE as:
$\beta^{g}(\tau)\ =\ Q^{g}_{Y(1)}(\tau)-Q^{g}_{Y(0)}(\tau),\quad\tau\in(0,1).$ (2)
where $Q^{g}_{Y(z)}(\cdot)$ is the quantile function of $F^{g}_{Y(z)}(y)$. Note, the introduction of $g(\cdot)$ permits the specification of alternative target populations to which the intent is to generalize. For example, the population QTE (Koenker and Bassett Jr, 1978; Koenker, 2005) is a special case of the WQTE for $g(x)=1$. As another example, if $g(x)$ equals the propensity score $e(x)=\textrm{pr}(Z=1\mid X=x)$, the estimand in equation (2) corresponds to the $\tau$th QTE among the treated, a QTE analogue to the average treatment effect among the treated. See Sun et al. (2021) for more details.
2.3 Identification and estimation given complete data
Identification of $\beta^{g}(\tau)$ is made possible under the following standard assumptions:
  1. 1.
    Stable unit treatment value assumption (SUTVA): For each unit, the potential outcomes are unrelated to the treatment status of other units (no interference), and there are no different forms or versions of each treatment level which could lead to different potential outcomes (consistency).
  2. 2.
    No unmeasured confounding: $\{Y(0),Y(1)\}\perp Z\mid X$.
  3. 3.
    Positivity: For some $c_{e}>0$, $c_{e}<e(x)<1-c_{e}$ for all $x\in\mathcal{X}$.
Focusing on the binary treatment setting, let $Q^{g}_{Y(0)}(\tau)\eqqcolon\beta^{g}_{0}(\tau)$ so that, by the definition given by expression (2), $Q^{g}_{Y(1)}(\tau)=\beta^{g}_{0}(\tau)+\beta^{g}(\tau)$. Sun et al. (2021) showed that under the three above assumptions, $Q_{Y(z)}^{g}(\tau)$ is identified by the following population-level estimating equation: for $z\in\{0,1\}$,
$0=E\left[g(X)\frac{\mathbbm{1}\left\{Z=z\right\}}{e(X)^{z}(1-e(X))^{1-z}}\left(\tau-\mathbbm{1}\{Y\leq Q_{Y(z)}^{g}(\tau)\}\right)\right].$
These estimating equations motivate the following estimator: given a sample of independent and identically distributed (i.i.d.) draws from the joint distribution $P_{c}$, say $\{(Y_{1},Z_{1},X_{1}),\ldots,(Y_{n},Z_{n},X_{n})\}$, Sun et al. (2021) proposed the IPW Z-estimator defined as the solution to:
$\mathbbm{P}_{n}\left(W^{g}(Z,X)(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}\right]\right)=0,\quad\tau\in(0,1),$ (3)
with vector $(1,Z)$ and
$W^{g}(Z_{i},X_{i})=\frac{g(X_{i})}{e(X_{i})}Z_{i}+\frac{g(X_{i})}{1-e(X_{i})}(1-Z_{i}).$ (4)
and where $\mathbbm{P}_{n}(U)=n^{-1}\sum_{i=1}^{n}U_{i}$ denotes the empirical mean of a generic random variable $U$.
3 The proposed framework
3.1 Double-sampling
For settings where $Y$ is potentially missing, let $R\in\{0,1\}$ denote the corresponding observance indicator so that $R=0/1$ if $Y$ is missing/observed. In this context, MAR states, intuitively, that the probability of $Y$ being observed only depends on the fully observed variables. Formally, one can write MAR as a conditional independence assumption: $R\perp Y\mid Z,X$. From this formal definition, MNAR may be seen to result when missingness depends on the underlying value of the outcome, either directly or through some (as-yet unspecified) other factor(s). In the bariatric surgery context, for example, the missingness process may be MNAR if patients who do not respond to surgery (i.e., fail to lose weight and/or maintain weight loss) are systematically more likely to have complete data—even after accounting for observed covariates and treatment status—perhaps because they are more likely to continue engaging with the health care system.
In settings where it is not clear whether MAR holds, we consider collecting the otherwise missing information on a sub-sample of patients, a process termed ‘double-sampling’ by Frangakis and Rubin (2001). Towards this, let $S\in\{0,1\}$ be an indicator for double-sampling, specifically indicating whether an individual with initially missing outcome data is selected into the follow-up sample and complete data are successfully collected. Note, by design, $S$ can only be $1$ if $R=0$ and is deterministically 00 if $R=1$; therefore, $S\equiv S(1-R)$. Additionally, the process by which these otherwise missing data are ascertained will be context specific and may involve active participation of study subjects (e.g., response to a survey or phone call) or no participation (e.g., manual chart review); see Levis et al. (2023) for additional details on practicalities. Finally, under this scheme, the ‘observed data’ are $n$ i.i.d. copies of $\{R,S,(R+S)Y,Z,X^{\top}\}$ with corresponding joint distribution $P_{o}$.
3.2 Assumptions
At the outset, we present a series of assumptions that will be leveraged to establish identification of the effects of interest, and the asymptotic properties of the estimator proposed in Section 3.3:
Assumption 1.
Non-informative double-sampling: $S\perp Y\mid Z,X,R=0$
Assumption 2.
Positivity of double-sampling probabilities: For some $c_{\eta}>0$, $c_{\eta}<\eta(z,x)\coloneqq\textrm{pr}(S=1\mid Z=z,X=x,R=0)<1-c_{\eta}$ for $z\in\mathcal{Z},x\in\mathcal{X}$.
Assumption 3.
For all $\tau\in(0,1)$, $z\in\mathcal{Z}$, at the true values $(\beta^{g*}_{0}(\tau),\beta^{g*}(\tau))$, the distribution function $F_{Y\mid Z=z,X}$ has one-sided derivatives at $\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)z$ with density $f_{Y\mid Z=z,X}$.
Assumption 4.
Letting $(\beta^{g*}_{0}(\tau),\beta^{g*}(\tau))$ denote the true values of $(\beta^{g}_{0}(\tau),\beta^{g}(\tau))$ for all $\tau\in(0,1)$, we have
  1. 1.
    $D_{1}(\tau)=E\left[g(X)f_{Y(1)\mid X}\{\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)\}\right]\in(0,\infty)$;
  2. 2.
    $D_{0}(\tau)=E\left[g(X)f_{Y(0)\mid X}\{\beta^{g*}_{0}(\tau)\}\right]\in(0,\infty)$.
Assumption 5.
Let $\eta(Z,X;\kappa)$ and $e(X;\gamma)$ denote user-specified parametric models for the double-sampling probabilities and the propensity scores, respectively. We have that $\eta(Z,X;\kappa)$ and $e(X;\gamma)$ are continuously differentiable at true parameters $\kappa^{*}$ and $\gamma^{*}$, respectively, with the derivatives being bounded uniformly in $X,Z$.
Assumptions 1 and 2 relate directly to the double-sampling scheme and are required to ensure that the full data distribution, $P_{f}$, is identified given the observed data. Specifically, Assumption 1 states that whether or not an individual is double-sampled is independent of the initially unobserved outcome, conditional on initially observed data. Meanwhile, Assumption 2 requires that among individuals with initially missing data, the conditional probability of being double-sampled cannot be 0 or 1 in any subgroup defined by $(Z,X)$.
Note, Assumption 1 can be viewed as an MAR assumption for the second phase of data collection. In practice, depending on the second phase sampling method, there may be a discrepancy between the intended sampling probabilities that will typically be under the control of the researchers, and the process by which data are actually collected (e.g., does an individual fill out the survey?). The follow-up sample, therefore, may or may not have complete data relative to the intended second stage sampling scheme. For example, when conducting a chart review without individuals’ direct engagement, it is reasonable to anticipate that those sub-sampled will provide complete data. In this setting, the plausibility of Assumption 1 may be quite likely. In other settings, however, such as when the data are collected via a telephone survey, some individuals initially selected by the design may decline to participate, leading to an additional layer of non-response (Koffman et al., 2021). In this situations, the double-sampling probabilities $\eta(Z,X)$ will have to be estimated, and the plausibility of Assumption 1 considered in light of this additional missingness.
Assumptions 3 and 4 are required to control the asymptotic behaviour of the proposed estimator. Briefly, Assumption 3 is standard in quantile-based analyses (Koenker and Bassett Jr, 1978; Koenker, 2005). Assumption 3 is necessary to ensure the expectation of the vector-valued estimating function ($\phi_{\theta}$ defined in expression (8)) is differentiable at the true values $\beta^{g*}_{0}(\tau)$ and $\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)$. Assumption 4 requires that the conditional density function be locally bounded such that the variance is finite, necessary to establish the asymptotic theorems in Section 3.4 and ‘Asymptotic theory for the entire quantile process’. We note that Assumption 4 is less restrictive than requiring the density function $f_{Y\mid Z,X}$ being continuous and uniformly bounded away from 0 and $\infty$ at $\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)Z_{i}$.
Finally, Assumption 5 clarifies assumptions regarding parametric specifications for the double-sampling probabilities and propensity scores when one or both require estimation. Specifically, implicit to the assumption is that the class of models adopted for each is sufficiently broad as to contain the true underlying mechanism. Practically, consideration of this assumption will be made in the context of whether and how flexibility is introduced into models specifications (e.g., with the use of splines). While, arguably, a focus on parametric models may be restrictive, it likely aligns with how most related clinical and public health research is conducted, and so represents an important case to consider. Nevertheless, possible extensions to nonparametric models are discussed Section 6.
3.3 An inverse-probability weighted estimator
In this work, we consider IPW-based estimators of population WQTEs (IPW-WQTEs). Specifically, supposing first that the propensity score and double sampling probabilities are known, we consider for any $\tau\in(0,1)$ the estimator defined as approximately solving the following estimating equation:
[…]
$E\left(\left\{R+\frac{S(1-R)}{\eta(Z,X)}\right\}W^{g}(Z,X)(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)Z\}\right]\right)=0$ (6)
at the true values $\{\beta^{g*}_{0}(\tau),\beta^{g*}(\tau)\}$, which motivates our approach. Finally, denote the solution to (5) as $\{\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$.
In practice, as discussed in Section  3.2, the true propensity score and, possibly, the double-sampling probability, will need to be estimated. Following Assumption 5, suppose the parameters indexing the models $\eta(Z,X;\kappa)$ and $e(X;\gamma)$ are estimated in-sample (i.e., using the data we have), obtained as the solutions to estimating equations based on estimating functions $\phi_{\kappa}$ and $\phi_{\gamma}$, respectively. Plugging in the estimated propensity scores and double-sampling probabilities into expression (5) yields the proposed estimator as approximately solving an equation:
$\mathbbm{P}_{n}\left(\left\{R+\frac{S(1-R)}{\eta(Z,X;\hat{\kappa})}\right\}{W^{g}}(Z,X;\hat{\gamma})(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}\right]\right)=0.$ (7)
Equivalently, letting $\theta=\{\kappa,\gamma,\beta^{g}_{0}(\tau),\beta^{g}(\tau)\}$ and $\Phi_{\theta}=(\phi_{\kappa},\phi_{\gamma},\phi_{\theta})$ with
$\phi_{\theta}=\left\{R+\frac{S(1-R)}{\eta(Z,X;\kappa)}\right\}W^{g}(Z,X;\gamma)(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}\right],$ (8)
[…]
3.4 Asymptotic theory for fixed quantiles
Suppose interest lies in estimating the causal WQTE for some fixed quantile level $\tau\in(0,1)$. The following result establishes pointwise properties of $\widehat{\beta}^{g}(\tau)$:
Theorem 1.
Under Assumptions 15, for any fixed $\tau\in(0,1)$,
$n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}=e_{2}^{\top}\Sigma^{-1}(\tau)\mathbbm{G}_{n}(\psi_{\theta^{*}})+o_{p}(1),$
where $e_{2}=(0,1)^{\top}$,
$\Sigma(\tau)=\begin{pmatrix}[l]D_{1}(\tau)+D_{0}(\tau)&D_{1}(\tau)\\ D_{1}(\tau)&D_{1}(\tau)\\ \end{pmatrix},$
$\psi_{\theta}=V_{\kappa,3}V^{-1}_{\kappa,1}\phi_{\kappa}+V_{\gamma,3}V^{-1}_{\gamma,2}\phi_{\gamma}-\phi_{\theta},$
$V_{\kappa,1}$ and $V_{\gamma,2}$ respectively are the derivative of $E(\phi_{\kappa})$ and $E(\phi_{\gamma})$, and $V_{\kappa,3}$ and $V_{\gamma,3}$ sequentially are the derivatives of $E(\phi_{\theta})$ at $\kappa$ and $\gamma$, and $\mathbbm{G}_{n}f=\sqrt{n}(\mathbbm{P}_{n}-P)f$.
Note that the vector $e_{2}$ is used to derive the linear representation for the proposed IPW-WQTE estimator $\widehat{\beta}^{g}(\tau)$, which is one of the two coefficients in equation (7). Assumption 4 ensures that the matrix $\Sigma(\tau)$ is invertible. Briefly, the proof of this theorem first involves showing that the class $\{\Phi_{\theta}:||\theta-\theta^{*}||<\delta\}$ is $P$-Donsker for some $\delta>0$ with finite envelope function. Theorem 19.26 of van der Vaart (2000) can then be invoked directly. The full proof of the above theorem is provided in Section S.2. of the Supplementary Materials. Crucially, Theorem 1 establishes that the proposed IPW-WQTE estimator, $\widehat{\beta}^{g}(\tau)$, can be written in a form of linear representation as a normalized sum of i.i.d. random variables, with a negligible remainder, so that, in turn, the central limit theorem can be applied:
Corollary 1.
Suppose that the assumptions in Theorem 1 hold. For any fixed $\tau\in(0,1)$, $n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}$ converges in distribution to $N(0,e_{2}^{\top}\Sigma^{-1}cov(\psi_{\theta^{*}})\Sigma^{-1}e_{2})$ as $n\rightarrow\infty$.
3.5 Asymptotic theory for the entire quantile process
For settings where interest lies in estimating causal WQTEs across a range of quantiles, we extend the asymptotic theory regarding $\widehat{\beta}^{g}(\tau)$ presented above to the entire process $\{\widehat{\beta}^{g}(\tau):\tau\in\mathcal{T}\}$ with $\mathcal{T}$ some compact subset of $(0,1)$. For an arbitrary set $\mathcal{T}$, let $\ell^{\infty}(\mathcal{T})$ denote the space of all bounded functions from set $\mathcal{T}$ to $\mathbbm{R}$ equipped with the uniform norm $\|f\|_{\mathcal{T}}=\sup_{\tau\in\mathcal{T}}|f(\tau)|$. Specifically, the following Theorem establishes key asymptotic properties of the process $M_{n}=\left\{n^{1/2}\{\widehat{\beta}^{g}(\tau)-\beta^{g*}(\tau)\}:\tau\in\mathcal{T}\right\}$:
Theorem 2.
Under the conditions of Theorem 1 hold, for any $\tau_{1},\tau_{2}\in\mathcal{T}$, the sequence of maps $M_{n}$ convergences weakly to a measurable tight process $M_{\infty}$ in $\ell^{\infty}(\mathcal{T})$ as $n\rightarrow\infty$, where $M_{\infty}$ is a centered Gaussian process with covariance
$cov\{M_{\infty}(\tau_{1})M_{\infty}(\tau_{2})\}=e^{\top}_{2}\Sigma^{-1}(\tau_{1})\Omega(\tau_{1},\tau_{2})\Sigma^{-1}(\tau_{2})e_{2}$
where $\Omega(\tau_{1},\tau_{2})$ is the asymptotic covariance between $\mathbbm{G}_{n}(\psi_{\theta^{*}})$ at $\tau_{1}$ and $\tau_{2}$.
See the Supplementary Materials for a proof.
3.6 Variance estimation via the bootstrap
Practically, variance estimation for both the fixed quantile and uniform inference settings will likely be challenging. In each case, the asymptotic variance has a complex form, and estimating it requires estimates of the potential outcome conditional densities (i.e., $f_{Y(1)\mid X}$ and $f_{Y(0)\mid X}$). Moreover, estimating these will be complex if, as is often the case, $X$ includes many covariates of different types. To simplify the variance calculation, we propose alternatives based on the bootstrap. For pointwise inference regarding some fixed quantile level, we note that any standard bootstrap approach can be used. For example, one could proceed by performing the same estimation procedure on a bootstrap replicate the same way as in the original data set. To construct uniform confidence bands and perform uniform inference on the entire IPW-WQTE process, we present the following theorem that justifies a procedure inspired by the gradient bootstrap resampling method proposed by Belloni et al. (2019). Our procedure differs, however, in that the coefficient process is derived from a series of weighted quantile regressions. Specifically, we define a gradient bootstrap estimator of $\beta^{g}(\tau)$ as a solution to the perturbed problem:
$\widehat{\beta}^{g^{*}}(\tau)\in\mathrm{argmin}_{\beta^{g}_{0}(\tau),\beta^{g}(\tau)}\mathbbm{P}_{n}$ [$\Bigg{[}\left\{R+\frac{S(1-R)}{\eta(Z,X;\hat{\kappa})}\right\}{W^{g}}(Z,X;\hat{\gamma})\rho_{\tau}\{Y-\beta^{g}_{0}(\tau)-\beta^{g}(\tau)Z\}$
$\quad-\mathbbm{U}^{*}(\beta^{g}_{0}(\tau),\beta^{g}(\tau);\tau)/\sqrt{n}\Bigg{]},$
where $\rho_{\tau}(u)=\tau\{\tau-\mathbbm{1}(u\leq\tau)\}$ is the check function, the parameters $\kappa,\gamma$ are estimated from the original data, and
$\mathbbm{U}^{*}(\beta^{g}_{0}(\tau),\beta(\tau);\tau)=\ \frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{R_{i}+\frac{S_{i}(1-R_{i})}{\eta(Z_{i},X_{i};\hat{\kappa})}\right\}{W^{g}}(Z_{i},X_{i};\hat{\gamma}){\{\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z_{i}\}}\{\tau-\mathbbm{1}(U_{i}\leq\tau)\},$
where $\{U_{i};i=1,\ldots,n\}$ is a sequence of i.i.d. $U(0,1)$ random variables. The following theorem justifies the validity of this gradient bootstrap estimator of $\beta^{g}(\tau)$ and thereafter its use to construct data-driven uniform confidence bands for $\beta^{g}(\tau),\tau\in\mathcal{T}$.
Theorem 3.
Suppose that the conditions in Theorem 1 hold. Then the sequence of maps $\hat{M}_{n}=\left[n^{1/2}\{\widehat{\beta}^{g}(\tau)-\widehat{\beta}^{g*}(\tau)\},\tau\in\mathcal{T}\right]$ convergences weakly to a measurable tight process $M_{\infty}$ in $\ell^{\infty}(\mathcal{T})$ defined in Theorem 2.
The proof of Theorem 3 can be found in the Supplementary Materials. As mentioned above, Theorem 3 allows us to compute (asymptotically) uniformly valid confidence bands for the process $\beta^{g}(\tau),\tau\in\mathcal{T}$ by simulating the process $\mathbbm{U}^{*}(\beta^{g}_{0}(\tau),\beta(\tau);\tau)$. The algorithm is provided in the Supplementary Materials.
In practice, if researchers do not have a predefined quantile level of interest and aim to estimate WQTEs across multiple quantile levels, Theorems 2 and 3 ensure asymptotically valid inference across an arbitrary compact set in $(0,1)$. Importantly, the uniform confidence bands are useful for making global inference across a compact set $\mathcal{T}$ rather than just a specific point; any null hypothesis that lies outside of the uniform confidence bands, even at a single quantile, can be rejected at a pre-specified significant level. For example, if the band fails to contain 0 at even one quantile in the set $\mathcal{T}$ of interest, we can reject the null hypothesis that there is no causal effect across $\mathcal{T}$.
4 Simulation
We conducted a series of simulation studies to investigate the properties of the methods proposed in Section 3. Throughout, for simplicity, we focus on estimation of the population QTE, that is $\beta^{g}(\tau)$ given by expression (2) with $g(X)=1$.
4.1 Data generation
For each simulation scenario we generated 10,000 datasets, each with sample size $n=10,000$. For each observation within each dataset, we generated two covariates: $X_{1}\sim U(0,1)$ and $X_{2}\sim U(0,2)$. Letting $X\equiv(X_{1},X_{2})$, a binary ‘treatment’ variable was generated as $Z\mid X\sim\mbox{Bernoulli}\{e(X)\}$, with $e(X)=\{1+\exp(-0.5+0.5X_{1}+0.5X_{2})\}^{-1}$. Finally, the outcome, $Y$, was generated according to the following linear regression model: $Y=1+Z+X_{1}+X_{2}+(1+\rho Z)\varepsilon$, where $\varepsilon$ is the error term and $\rho$ is a constant controlling the heteroscedasticity of QTEs across quantiles $\tau\in(0,1)$. For the error terms, we considered a Pareto distribution with location parameter 1 and shape parameter 5. With this choice, the induced marginal distribution of outcome $Y$ is highly right skewed, as shown in Figure S.1 in the Supplementary Material. For $\rho$, we considered two scenarios: (i) $\rho=0$, so that the QTE is homogeneous in across quantile levels; and, (ii) $\rho=1.5$, so that the QTEs are heterogeneous across quantile levels. The true QTE is 1 for the homogeneous setting, and the true QTEs vary from 2.67 to 3.10 for the heterogeneous setting; see Figure S.2 in the Supplementary Material.
Missingness was subsequently introduced to reflect the primary scenario that the proposed methods were designed to address, that being missing outcomes subject to MNAR. Towards this, we generated an initial observance indicator via, $R\mid Y\sim\mbox{Bernoulli}\{\pi(Y)\}$: for the homogeneous QTE scenario, we set $\pi(Y)=(1+\exp(-1-4.3Y+Y^{2}))^{-1}$; for the heterogeneous QTEs scenario, we set $\pi(Y)=(1+\exp(-1-3.8Y+Y^{1.8}))^{-1}$. Note, both of these specifications resulted in marginal missingness probabilities of approximately 35%.
Finally, we considered double-sampling of those with $R=0$ within strata defined jointly by $(Z,X_{1},X_{2})$, with $X_{1}$ and $X_{2}$ dichotomized using cut-off points of 0.5 and 1, respectively. We pre-defined the stratum-specific sample size varying from 0 and 240 in each of the eight strata. The specific sample size for each stratum is listed in Section S4 of the Supplementary Materials. To ensure the non-informative double-sampling assumption, a stratum-specific sample was randomly ‘selected’ in each stratum. This procedure ‘selected’ a total of double-sampled 800 individuals with initially missing outcomes. Overall these choices resulted in double-sampling of approximately 22% of individuals with original missing data.
4.2 Analysis
We consider estimation of the population QTE, $\beta^{g}(\tau)$ with $g(X)=1$, with values of $\tau$ ranging from 0.1 to 0.9 with a grid spacing of 0.1. Specifically, we considered four estimators: (I) an estimator based on the full data, with $e(X)$ estimated via maximum likelihood; (II) an estimator based on the initial incomplete data alone (i.e., only including those with complete outcome data at the outset and not including the double-sampled data), with $e(X)$ estimated via maximum likelihood; (III) an estimator based on the initial incomplete data coupled with the double-sampled data, with $e(X)$ set at their true values and $\eta(Z,X)$ estimated via maximum likelihood; and, (IV) an estimator based on the initial incomplete data coupled with the double-sampled data, with both $e(X)$ and $\eta(Z,X)$ estimated via maximum likelihood. Note, throughout, estimates of $e(X)$ were based on a correctly-specific logistic regression. Furthermore, since we performed a stratified sampling with fixed sample sizes, and the eight potential strata had random sample sizes (due to the re-generation of the covariate data), there is no single set of ‘true’ values for the double-sampling probabilities. We thus estimated $\eta(Z,X)$ via a saturated logistic regression on the eight strata.
Asymptotic variance calculations for proposed estimators (III) and (IV) were computed assuming the true underlying distribution family. Specifically, for estimator (III), the function $\psi_{\theta}$ from Theorem 1 simplifies to $V_{\kappa,3}V^{-1}_{\kappa,1}\phi_{\kappa}-\phi_{\theta}$. The asymptotic variance was then calculated with the true $e(x)$ and estimated parameters $\{\widehat{\kappa},\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$ plugged into $e_{2}^{\top}\Sigma^{-1}cov(\psi_{\theta})\Sigma^{-1}e_{2}$. For estimator (IV), we first estimated parameters $\widehat{\theta}=\{\widehat{\kappa},\widehat{\gamma},\widehat{\beta}_{0}^{g}(\tau),\widehat{\beta}^{g}(\tau)\}$, then plugged them into the corresponding asymptotic variance expression of (IV), as presented in Theorem 1 and Corollary 1. Additionally, pointwise bootstrap variance estimates were computed for estimators (III) and (IV), for each $\tau$ considered. Finally, we constructed 95% confidence bands using the gradient bootstrap method described in Section 3.5. To assess coverage for the latter, we determined whether a given confidence band contained the true QTE value for each quantile level in the sequence considered. If the band contained all of the true values, we consider it successful coverage for the given realization; otherwise, it is considered a failure. The empirical coverage probability is the proportion of realizations with successful coverage of the true parameters.
4.3 Results
Figure 1 presents point-wise relative bias ($\times$ 100) and 95% confidence interval coverage for the homogeneous and heterogeneous QTE scenarios. Throughout, we see that the proposed estimators (III) and (IV) in Section 4.2 have relative bias and confidence interval coverage comparable to the estimator (I) based on the full data. Regarding the 95% confidence bands for uniform inference, coverage was 96.7% for estimation of the homogeneous QTE and 98.0% for estimation of the heterogeneous QTE.
Finally, Figure 2 illustrates how results might be presented in practice, that is for a single dataset. In particular, Figure 2 compares 95% pointwise confidence intervals, including both the asymptotic variance formula-based and bootstrap-based intervals, versus the confidence bands for a realization, both for homogeneous and heterogeneous QTE settings. The black dotted line shows the point QTE estimates, the blue and yellow dashed lines show the 95% pointwise confidence intervals, and the light grey areas show the 95% uniform confidence bands, which are wider by construction. In the left panel of Figure 2, under a homogeneous QTE setting, the QTE estimates demonstrate similarity across various quantile levels; whereas in the right panel, the QTE estimates show more beneficial treatment effects with larger $Y$ values.
5 Data Application
In this section, we illustrate proposed methods proposed with a comparison of the relative effects of two bariatric surgery procedures on long-term weight loss using EHR data from a broader investigation of long-term outcomes for patients who had undergone bariatric surgery in one of the three Kaiser Permanente health care sites: Washington, Northern California, and Southern California(Arterburn et al., 2021). At each study site, information on health system enrollment, diagnoses, procedures, medical records, laboratory values, and other data systems were routinely recorded for research.
In the present study, we aim to compare Roux-en-Y gastric bypass (RYGB) ($Z=1$) versus vertical sleeve gastrectomy (VSG) ($Z=0$) bariatric surgery procedures on percent BMI change at 3 years post-surgery ($Y$). From a larger sample of 30,991 patients for which follow-up outcomes were partially observed, we selected $13,514$ patients who underwent RYGB or VSG between January 2005 and September 2015, with complete weight data at baseline (closest measurement pre-surgery, up to 6 months) and at follow-up (closest measurement within $\pm$ 90 days). See Table S.3 in the Supplementary Materials for a summary of the baseline characteristics of individuals who underwent bariatric surgery procedures.
For the study sample of $13,514$ patients, we then artificially imposed a MNAR missingness mechanism in the outcome. To accomplish this, we first modelled the probability of observance indicator ($R$) from the original larger population of 30,991 patients using the following baseline covariates ($X$): baseline weight, health care site, year of surgery, age, gender, ethnicity, number of days of health care use in 7-12 month period pre-surgery, number of days hospitalized in pre-surgery year, smoking status, Charlson/Elixhauser comorbidity score, insurance type, clinical statuses for hypertension, coronary artery disease, diabetes, dyslipidemia, retinopathy, neuropathy, mental health disorders, and use of medicines which includes insulin, ACE inhibitors, ARB, statins, other lipid lowering medications, and other antihypertensives. The model for $\textrm{pr}(R=1\mid Z,X)$ was fit using a SuperLearner ensemble with library {SL.glm, SL.ranger, SL.rpart} in R version 4.2.1 (Polley et al., 2023).
Subsequently, to impose MNAR, we additionally included a dependence on the outcome $Y$ by augmenting the fitted values $\widetilde{p}$ of $\textrm{pr}(R=1\mid Z,X)$ via:
[…]
where ${Y^{*}}$ is the standardized BMI change at 3 years. Finally, we trimmed $p$ using values $0.05$ and $0.95$.
To generate MNAR outcomes on the sample of 13,514 patients, we sampled $R$ according to a Bernoulli with probability given by the trimmed $p$. The resulting marginal probability of missingness was 30%. We considered double sampling random subsamples of size 1,000 and 1,500 from all patients with initially missing outcomes. For each of the two resulting data sets, we computed the QTEs at a sequence of quantile levels (i.e., $0.1,0.3,0.5,0.7,0.9$), and compared the point estimates and 95% pointwise bootstrap confidence intervals using the proposed method in Section 3. A benchmark QTE estimate for each quantile level was estimated using the full data set (before artificial imposition of missing outcomes). For comparison, we also estimated QTEs using the complete-case data set ($R=1$). We computed 95% confidence bands using the data set with 1,500 double-sampled patients and compared to the pointwise bootstrap confidence intervals. Additionally, as the effects of the two bariatric surgeries may be different among patient with diabetes or dyslipidemia diagnosis, we repeated the above analyses stratifying by diabetes and dyslipidemia diagnosis categories.
The left column of Figure 3 displays the pointwise QTE estimates and 95% confidence intervals across the quantile levels, while the right column compares pointwise CIs to confidence bands constructed using the procedure in Section 3.6. Panels (A) to (C) of Figure 3 represent different subsets of the data: (A) are for the overall 13,514 patients, while row (B) and (C) are stratified by diabetes status and dyslipidemia diagnosis, respectively. The specific estimated QTE values and 95% pointwise CIs are listed in Table S.4 in the Supplementary Materials.
From the results we see that the complete case estimates are biased when compared to the benchmark estimates that use the full data, especially at low quantile levels. The proposed IPW-WTQE estimators appeared to be unbiased, with the setting with 1,500 subjects in the second-stage sample showed greater efficiency (i.e. with narrower 95% CIs) compared to the setting with the smaller subsample. Heterogeneous QTEs were observed across quantile levels as well as at different strata: when compared to VSG, RYGB appears to be associated with greater percent BMI loss at 3 years: (i) at low tails of the percent BMI change distribution, and (ii) among patients with diabetes or dyslipidemia diagnosis at most of the selected quantile levels.
6 Discussion
This paper proposes a novel approach for estimation and inference for causal weighted QTEs in settings where outcome data are potentially MNAR. The approach combines additional data collection on a sub-sample of study units with otherwise missing information, with a set of proposed identifying assumptions. When estimating contrasts that involve the mean of the counterfactual distribution such as the ATE, one may employ $g$-methods that first compute a conditional ATE, conditional on $X$, and then marginalize to recover an estimate of the population causal effect (Robins, 1986; Hernán and Robins, 2020). However, due to Jensen’s inequality, integrating a conditional QTE will not yield the population QTE (Sun et al., 2021). In this work, therefore, we developed a novel inverse-probability weighted estimator as well as a bootstrap-based approach for valid inference pointwise for a single quantile or uniformly for a range of quantiles. The procedure involves first estimating propensity scores and double-sampling probabilities, and then plugging these into an estimating function for the weighted QTE. In a simulation study, we consider settings where the QTEs are either homogeneous or heterogeneous across quantiles, verify theoretical estimation properties in a finite sample setting and demonstrate validity of the proposed bootstrap-based confidence intervals/bands.
Key to the proposed methods is that no assumptions are required regarding the missingness mechanism in the original data (i.e., for $R$). This, arguably, is a critical benefit in that the mechanism(s) underlying $R$—especially in the increasingly common setting in which EHR data are being used—are typically outside of the control of the research team and potentially quite complex. As such, even thought experiments regarding assumptions required to make progress with the original data (e.g., to consider the potential validity of MAR) can be challenging and inconclusive. One way of framing the proposed methods, therefore, is that the need to make assumptions regarding $R$ (as would have to be the case if only the original data were available) are replaced with the need to make assumptions regarding the double-sampling indicator $S$; that is, Assumptions 1 and 2. In practice, depending on the context and the nature of the additional data to be collected, consideration of these assumptions may be fairly straightforward. For example, suppose the additional data collection involves manual chart review or natural language processing of a sub-sample chosen by the research team (Weiskopf et al., 2019). In this setting, since no involvement of the study subject is needed, Assumptions 1 and 2 may hold trivially by design. In other settings, however, in particular when participation of a study subject is needed (e.g., to respond to a survey or phone call), assessment of these assumptions may require more detailed consideration. However, even in these situations, researchers may pursue additional design options that may shed light. For example, researchers may employ intensive, repeated efforts to contact study subjects and collect the data and/or engage with study subjects who have complete information in the original data to learn about potential recall bias (Haneuse et al., 2016; Koffman et al., 2021). EHR studies often encompass vast datasets, yet the ability to reach out to patients typically does not grow with sample size. Given that the double-sampling framework can be strategically planned from the beginning, it presents an opportunity to allocate resources effectively within the budgetary confines. Our ongoing research aims to find an optimal sampling rule that minimizes the semiparametric efficiency bound subject to a budget constraint.
There are numerous opportunities for building on the proposed framework. One such opportunity is to consider settings beyond binary treatment settings, including more than two discrete options as well as continuous treatments (e.g., dose) (Sun et al., 2021). A second opportunity is to consider alternatives to parametric models for the propensity scores and double-sampling probabilities, such as more flexible machine learning techniques (Chernozhukov et al., 2018). For the latter, an important challenge will be that the rate of convergence for estimation in non-parametric models will affect both the achievable rates of convergence and asymptotic distribution of the weighted QTE estimator. Relatedly, we note that Belloni et al. (2019) developed a theory for inferring the unweighted QTE in the absence of missing data, under specific assumptions on the nonparametric models for the quantile regression function. Even in this case without missing data, however, the complete picture of nonparametric estimation of the unweighted QTE through the lens of a higher-order semiparametric theory (Robins et al., 2008) remains open. Therefore, here we keep a detailed study of the nonparametric analogue of the problem we consider for future directions. Finally, it will often be the case that covariates beyond the outcome may be missing (e.g., an important confounder); extending the assumptions and estimator to acknowledge this is an avenue we are currently investigating.
7 Supplementary Materials
Supplementary material available online includes the proofs of the results in the main manuscript, the gradient bootstrap computational algorithm, and additional tables and figures.
8 Competing interests
No competing interest is declared.
9 Acknowledgements
The authors gratefully acknowledge support from National Institutes of Health grants R-01 DK128150 and R-01 HL166324.